# A Hybrid TDMA/CSMA Protocol for Time-Sensitive Traffic in Robot Applications

### v1

@@ Abstract
Recent progress in robotics has underscored the demand for real-time control in applications such as manufacturing, healthcare, and autonomous systems, where the timely delivery of mission‑critical commands under heterogeneous robotic traffic is paramount for operational efficacy and safety. In these scenarios, mission‑critical traffic follows a strict deadline‑constrained communication pattern: commands must arrive within defined QoS deadlines, otherwise late arrivals can degrade performance or destabilize control loops. In this work, we demonstrate on a real‑time SDR platform that CSMA, widely adopted in robotic communications, suffers severe degradation under high robot traffic loads, with contention‑induced collisions and delays disrupting the on‑time arrival of mission‑critical packets. To address this problem, we propose an IEEE 802.11‑compatible hybrid TDMA/CSMA protocol that combines TDMA’s deterministic slot scheduling with CSMA’s adaptability for heterogeneous robot traffic. The protocol achieves collision-free, low-latency mission-critical command delivery and IEEE 802.11 compatibility through the synergistic integration of sub-microsecond PTP-based slot synchronization—essential for establishing precise timing for TDMA, a three-session superframe with dynamic TDMA allocation for structured and adaptable traffic management, and beacon-NAV protection to preemptively secure these critical communication sessions from interference. Emulation experiments on real-time SDR testbed and Robot Operating System (ROS) simulation show that the proposed protocol reduces missed‑deadline errors by 93% compared to the CSMA baseline. In high‑speed robot path‑tracking ROS simulations, the protocol lowers Root Mean Square (RMS) trajectory error by up to 90% compared with a CSMA baseline, all while maintaining throughput for non‑critical traffic within ±2%.
Index Terms:
Real-time robot control, hybrid TDMA/CSMA protocol, command traffic congestion, deadline-constrained communication
@@ I Introduction
As the number of robots grows rapidly in a multi-robot system, communications between robots are becoming increasingly data-intensive. The massive amount of information exchanged across different devices — ranging from sensor data update to synchronized actuation — shares the same communication medium, making efficient multiple access a significant challenge for the timely and reliable delivery of mission-critical control commands.
@@
However, as the number of cooperating robots and their traffic loads increase, CSMA suffers from severe contentions between data streams, resulting in elevated latency when traffic congestion happens due to transmission back-offs [12, 13]. In multi‑robot cooperation, contention between routine data and time‑critical commands introduces delays that compromise the responsiveness required for real‑time tasks such as path following and formation control. Moreover, multi‑robot teams are inherently dynamic: robots may join or leave, links fluctuate, and task priorities shift. In such environments, static or slow‑to‑adapt access schemes quickly fall out of synchronization with demand, wasting resources on inactive nodes while starving newly critical flows [14]. Practical deployments must also scale to multiple robots and mixed‑traffic loads without collapsing best‑effort throughput; otherwise, safeguarding time-critical commands risks undermining large‑volume data and event‑driven messages essential for long‑term missions.
@@
Therefore, a practical robotic Medium Access Control (MAC) protocol must meet three concurrent challenges: (i) strict latency guarantees for time-sensitive command traffic under variable contention, (ii) rapid reconfiguration under dynamic topology and heterogeneous loads, and (iii) scalability in mixed‑traffic intensity without collapsing non‑time-critical long-term throughput.
@@
Prior robotic MAC protocols largely relied on either pure CSMA or pure TDMA, yet both are inadequate for real-time multi‑robot workloads. In [15], the authors adopt the non‑beacon enabled IEEE 802.15.4 CSMA/CA mode to keep nodes continuously active and deliver broadcasts immediately without waiting for coordinator beacons or slot schedules. However, CSMA still suffers from backoffs, hidden terminals, and collision bursts as team size or interference increases, leading to unbounded latency and jitter that threaten real‑time deadlines. Pure TDMA‑based designs, such as UMCI‑MAC [16] for underwater cooperation and RoboMAC [17] for low‑rate mobile teams, avoid collisions and bound delay through deterministic slot scheduling, but have key limitations. Their rigid cycle structure makes it difficult to support heterogeneous and dynamic robot traffic: urgent or bursty flows must often wait for the next cycle, while best‑effort traffic is inefficiently handled. Moreover, strict slot synchronization adds overhead, and as team size grows, longer cycles increase latency and constrain scalability.
Consequently, there is a pressing need for a specially designed communication protocol that ensures deterministic delivery of time-sensitive traffic while accommodating the diverse and dynamic demands of latency-insensitive traffic in multi-robot systems [10].
@@
Researchers have explored hybrid TDMA/CSMA protocols that combine 1) the deterministic scheduling in TDMA with 2) the flexibility of random access in CSMA. However, these designs were developed for generic wireless networks or other applications rather than robotic applications: they typically assume static or slowly varying conditions, prioritize average throughput or energy efficiency while overlooking strict deadlines. These limitations make them unsuitable for hard‑deadline command loops in multi‑robot cooperation over shared 802.11 links.
@@
In [18], the hybrid TDMA/CSMA protocol targets a fixed single‑hop star with long coordinator‑scheduled superframes, where TDMA handshakes and “slot borrowing” only take effect in the next superframe—leading to slow reallocation and weak support for frequent join/leave or urgent preemption. Synchronization is only beacon‑level (microsecond‑scale) and there is no session‑level protection, making it unsuitable for dynamic robot teams requiring rapid priority shifts. In [19], the MAC splits the contention access period into TDMA/CSMA under IEEE 802.15.4 standard and achieves energy and throughput gains at low rates, but its low data rate and long backoffs are incompatible with hard‑deadline robot control loops. In [20], hybrid access is formulated via Markov Decision Process: centralized schemes require global queue knowledge and heavy computation, while distributed schemes are vulnerable to congestion and hidden terminals, both hindering responsiveness and scalability as nodes or traffic patterns shift. In [21], soft‑TDMA divides associated nodes into coarse time‑frames with intra‑group CSMA. However, due to contention and only loose frame‑level synchronization, it offers no hard delay bounds. The scheme also provides weak isolation, leaving strict command‑deadline guarantees unmet under mixed robotic traffic.
@@
3. Experimental Validation: We validate our framework through SDR-based signal transmission experiments under realistic multi-robot traffic conditions. We use the resulting data for emulation experiments in a ROS environment employing the PurePursuit path-tracking algorithm for performance assessment. Experimental results demonstrate the system’s significant improvements in latency, packet loss, and throughput compared to traditional CSMA systems in multi-robot scenarios. See Section IV for details.
@@ II-A Data Traffic in Robotic Networks
For example, the integration of multimodal large language models (MLLMs) has enabled robots to understand human vocal input and generate appropriate responses to fulfill user requirements. The task-fulfillment process, as illustrated in recently proposed LLM-powered robotics and automation frameworks [24, 25], typically involves complex interactions between humans and multiple robots. For instance, in the LLMind framework [26, 27], a master robot receives human commands and sends polling messages to its slave robots to query their states. Each slave robot then reports its current status (e.g., availability, position, computational capability) back to the master for decision-making, which is followed by task decomposition at the master node and subtask execution at the slave devices. The above communication process is sporadic and bursty, given that: 1) the whole process is triggered by external, unpredictable human inputs, and 2) the response time of LLM-supported robots  [^1] In LLMind-like systems, both master and slave robots in are empowered by LLM agents. is variable due to fluctuating task complexity and device states. The event-driven traffic in this scenario is also less sensitive to communication latency. Furthermore, the primary source of delay is the LLM inference process on robots, which may take several tens of seconds even with lightweight embedded LLMs.
@@
The third class, mission-critical control traffic, refers to time-sensitive communication flows in multi-robot systems for control purposes. Unlike time-insensitive multi-robot tasks, such as those demonstrated with the LLMind framework [26, 27], in which a high-level subtask description is assigned to a device and get locally executed under the control of the embedded AI in the robot, a time-critical control command is generated by a central controller or a master robot, and it contains specific control messages that directly regulate the action of a slave robot. Typical examples of this type of traffic include the control of high-speed motors or real-time trajectory updates for distributed robots. As the control of robotic operations requires ultra-high precision, the control message must be delivered timely to ensure the operation is accurate and updated - even microsecond-scale delays can destabilize control loops, leading to potential system failures [10].
To support the timely distribution of mission-critical control messages from the master robot to slave robots, ROS has adopted a Publish-Subscribe (Pub/Sub) architecture built atop the Data Distribution Service (DDS) middleware. In the Pub/Sub architecture, the master node publishes control messages (e.g., velocity commands) to designated topics, typically using the Real-Time Publish-Subscribe (RTPS) protocol over TCP. Meanwhile, communication requirements are specified via Quality of Service (QoS) parameters within DDS, among which the QoS DEADLINE policy is pivotal for real-time robot controls — it specifies the maximum allowable interval between consecutive messages to enforce stringent timing requirements. For example, a 100 ms deadline requires the publisher to send an update at least every 100 ms, and the subscriber (e.g., an actuator controller that executes the command) expects to receive messages within this timing window. If this timing constraint is violated — probably due to delays in message generation or network transmission — the DDS middleware triggers events to alert the system. In such cases, the delayed command is no longer useful for execution, since the control action it represents has already expired.
@@ II-B Motivation Demonstrated via Emulation Experiments
Figure 3: A photo of the real-world system. setup
@@
3) The mission-critical robot command traffic was represented by a deadline-constrained packet in ROS, with the QoS deadline request set to 100ms. That communication link was realized by the Openwifi AP and the Openwifi STA222 The mission-critical robot command traffic can also be simulated with commercial AP and STAs. Here we use the more complex openwifi platform to realize that traffic link to provide a fair comparison for our later experiment: our later implementation of the hybrid TDMA/CSMA system was built upon the openwifi system, simulating the pure CSMA case with the same hardware platform would provide a fair baseline for later comparisons..
@@
Traffic Type Description
@@
Event-driven traffic An LLMind system with the following setup: 1. Server to Client: 2 KB (Polling) 2.Client to Server: 512 KB (Historical Data)3 3. Server to Client: 5 KB (FSM Data)4
@@
  • 4
    FSM Data is the finite-state machine representation of the decomposed task generated by the server to guide subsequent control script execution in LLMind system.
@@ III-A Design Principles of the Hybrid TDMA/CSMA Protocol
Third, for multi-robot cooperation, potential collisions may occur between TDMA and CSMA sessions due to their different access mechanisms. CSMA’s contention‑based channel access may overlap with TDMA slots or the control slot of TDMA, causing interference under diverse traffic conditions. Therefore, an effective mechanism is required to guarantee the deterministic delivery of TDMA packets while efficiently using residual bandwidth for CSMA to manage dynamic, non-critical robot traffic.
To address these challenges, our hybrid TDMA/CSMA protocol has three key features. First, we integrate the Precision Time Protocol (PTP) for sub-microsecond synchronization across nodes to support IEEE 802.11-compatible precise alignment of TDMA slots despite robot mobility. Second, we introduce a three-session superframe that separates traffic into: (i) deterministic TDMA slots for mission-critical control traffic; (ii) a TDMA control session for transmitting dynamic slot allocation information, and (iii) a general-purpose CSMA session for non-critical robot traffic. Finally, we employ the Network Allocation Vector (NAV) mechanism, embedded in beacon frames, to protect both the TDMA session and TDMA control information by preemptively reserving the channel. This prevents interference from opportunistic CSMA accesses and guarantees the deterministic, low-latency communication essential for multi-robot cooperation.
@@ III-B1 Feature 1: PTP Synchronization Mechanism
Precise time synchronization is essential for aligning time slots in the TDMA session. We achieve a sub-microsecond clock alignment across nodes through a three-way handshake mechanism, compensating for clock offsets and propagation delays.
@@
$\tilde{s}_{i,beacon}^{l}=s_{ap,beacon}^{l}+{{d}_{i,delay}}+{{o}_{i}}$
For synchronization purposes, any one of the slave clients should respond to AP with another synchronization message (referred to as a synchronization response message in the rest of this paper) at any time during session 2. We require each client accepted into the network to perform the synchronization procedure periodically. Specifically, for the synchronization process, the client embeds its transmission time, $s_{i,response}^{l}$, in the packet’s payload during the l-th superframe, representing the time the packet is sent according to the client’s local clock. The time $s_{i,response}^{l}$ is determined by the client’s local scheduling, initially based on the beacon’s arrival time, and progressively adjusted as the synchronization process continues (see Sections IV-A3 and IV-C in [22]). The server records the packet’s arrival time, $t_{i}^{l}$, and includes it in the next beacon, which is received by the client. Using these timestamps and assuming symmetric propagation delays ($d_{i}=d_{i,0}$), clients estimate:
${{d}_{i}}=\frac{1}{2}(\tilde{s}_{i,beacon}^{l}+t_{i}^{l}-s_{ap,beacon}^{l}-s_{i,response}^{l})$
${{o}_{i}}=\frac{1}{2}(\tilde{s}_{i,beacon}^{l}-t_{i}^{l}-s_{ap,beacon}^{l}+s_{i,response}^{l})$
To align packet arrivals at the server, the $i$-th clients schedules its transmission time for the $j$-th slot of the $l$-th superframe at Session 1 as follows:
@@
III-B2 Feature 2: Multi-Session Superframe with Flexible Time Slot Allocation

### v2

@@ Abstract
Recent progress in robotics has underscored the demand for real-time control in applications such as manufacturing and healthcare systems, where the timely delivery of mission‑critical commands under heterogeneous robotic traffic is paramount for operational efficacy and safety. In these scenarios, mission‑critical traffic follows a strict deadline‑constrained communication pattern: commands must arrive within defined deadlines, otherwise late arrivals can degrade performance or destabilize control loops. In this work, we demonstrate on a real‑time software-defined radio (SDR) platform that CSMA, widely adopted in robotic communications, suffers severe degradation with contention‑induced collisions and delays disrupting the on‑time arrival of mission‑critical packets. This degradation arises under a common robotic traffic pattern where non‑critical traffic dominates the channel, while lightweight mission‑critical commands must be delivered frequently with strict deadlines over the shared medium. To address this, we propose an IEEE 802.11‑compatible hybrid TDMA/CSMA protocol that combines TDMA’s deterministic slot scheduling with CSMA’s adaptability for heterogeneous robot traffic. The protocol achieves collision-free, low-latency mission-critical command delivery and IEEE 802.11 compatibility through the synergistic integration of sub-microsecond PTP-based slot synchronization, a three-section superframe with dynamic TDMA allocation for structured and adaptable traffic management, and beacon-NAV protection to preemptively secure critical communication applications from interference. Emulation experiments on a real-time SDR testbed show that the proposed protocol reduces missed‑deadline errors by 93% compared to the CSMA baseline under a robotic traffic setup at an overall aggregate channel load of 77.1%, wherein 99.9% of the traffic is from non time-critical applications and 0.1% of the traffic is from deadline-constraint applications. In a high‑speed robot path‑tracking Robot Operating System (ROS) simulation, the protocol lowers root mean square trajectory error by up to 90% compared with the CSMA baseline, while maintaining throughput for non‑critical traffic within ±2%.
@@ I Introduction
As the number of robots grows rapidly in a multi-robot system, communications between robots are becoming increasingly data-intensive. The massive amount of information exchanged across different devices – ranging from sensor data update to synchronized actuation – shares the same communication medium, making efficient multiple access a significant challenge for the timely and reliable delivery of mission-critical control commands.
@@
However, as the number of cooperating robots and their traffic loads increase, CSMA suffers from severe contentions between data streams, resulting in elevated latency when traffic congestion happens due to transmission back-offs [12, 13]. In multi‑robot cooperation, contention between routine data and time‑critical commands introduces delays that compromise the responsiveness required for real‑time tasks such as path following and formation control. Moreover, multi‑robot teams are inherently dynamic: robots may join or leave, links usages fluctuate, and task priorities shift. In such environments, static or slow‑to‑adapt access schemes quickly fall out of synchronization with demand, wasting resources on inactive nodes while starving newly critical flows [14]. Practical deployments must also scale to multiple robots and mixed‑traffic loads without collapsing best‑effort throughput; otherwise, safeguarding time-critical commands may undermine large‑volume data and event‑driven messages unnecessarily.
@@
A practical robotic Medium Access Control (MAC) protocol must meet three concurrent challenges: (i) strict latency guarantees for time-sensitive command traffic under variable contention, (ii) rapid reconfiguration under dynamic topology and heterogeneous loads, and (iii) scalability in mixed‑traffic intensity without collapsing non‑time-critical long-term throughput.
@@
Prior robotic MAC protocols largely relied on either pure CSMA or pure TDMA, yet both are inadequate for real-time multi‑robot workloads. In [15], the authors adopted the non‑beacon enabled IEEE 802.15.4 CSMA/CA mode to keep nodes continuously active and deliver broadcasts immediately without waiting for coordinator beacons or slot schedules. However, CSMA may suffer from backoffs, hidden terminals, and collision bursts as number of communicating robots increases, leading to unbounded latency and jitter that threaten real‑time deadlines. Pure TDMA‑based designs, such as UMCI‑MAC [16] for underwater cooperation and RoboMAC [17] for low‑rate mobile robotic teams, avoid collisions and bound delay through deterministic slot scheduling, but have their own limitations. Their rigid cycle structure makes it difficult to support heterogeneous and dynamic robot traffic: urgent or bursty flows must often wait for the next cycle, while best‑effort traffic is inefficiently handled. There is a need for a specially designed communication protocol that ensures deterministic delivery of time-sensitive traffic while accommodating the diverse and dynamic demands of latency-insensitive traffic in multi-robot systems.
@@
Researchers have explored hybrid TDMA/CSMA protocols that combine 1) deterministic scheduling in TDMA with 2) flexibility of random access in CSMA. However, these designs were developed for generic wireless networks or other applications rather than robotic applications. They typically assume static or slowly varying conditions, prioritizing average throughput while overlooking strict deadlines. These limitations make them unsuitable for hard‑deadline command loops in multi‑robot cooperation over shared mediums.
@@
The hybrid TDMA/CSMA protocol in [18] targeted a fixed single‑hop star with long coordinator‑scheduled superframes, where TDMA handshakes and “slot borrowing” only take effect in the next superframe – leading to slow reallocation and weak support for frequent join/leave or urgent preemption. In [19], the MAC splits the contention access period into TDMA/CSMA under IEEE 802.15.4 standard and achieves energy and throughput gains at low rates, but its low data rate and long backoffs are incompatible with hard‑deadline robot control loops. In [20], soft‑TDMA divides associated nodes into coarse time‑frames with intra‑group CSMA. However, due to contention and only loose frame‑level synchronization, it offers no hard delay bounds. The scheme also provides weak isolation, leaving strict command‑deadline guarantees unmet under mixed robotic traffic.
@@
3. Experimental Validation: We validate our framework through SDR-based signal transmission experiments under realistic multi-robot traffic conditions. We use the resulting data for emulation experiments in a ROS environment employing the PurePursuit path-tracking algorithm for performance assessment. Experimental results demonstrate that the system achieves deterministic, low-latency delivery of mission-critical commands without degrading the throughput of best-effort, non-critical traffic compared to traditional CSMA in multi-robot scenarios. See Section IV for details.
@@ II-A Data Traffic in Robotic Networks
In practical robotic networks, non‑critical robot traffic often dominates throughput, while lightweight mission‑critical commands must be delivered frequently under strict deadlines. In collaborative manufacturing[6], camera streams and logs occupy most bandwidth, yet small control commands such as velocity or pose updates demand tight timing guarantees. A similar pattern arises in surgical robots[14], where high‑definition video and sensor traffic prevail, but small critical commands like motion corrections or micro‑positioning updates face the strictest latency constraints to avoid risks.
@@
For example, the integration of multimodal large language models (MLLMs) has enabled robots to understand human vocal input and generate appropriate responses to fulfill user requirements. The task-fulfillment process, as illustrated in recently proposed LLM-powered robotics and automation frameworks [23, 24], typically involves complex interactions between humans and multiple robots. For instance, in the LLMind framework [25, 26], a master robot receives human commands and sends polling messages to its slave robots to query their states. Each slave robot then reports its current status (e.g., availability, position, computational capability) back to the master for decision-making, which is followed by task decomposition at the master node and subtask execution at the slave devices. The above communication process is sporadic and bursty, given that: 1) the whole process is triggered by external, unpredictable human inputs, and 2) the response time of LLM-supported robots [^1] Both master and slave robots in LLMind-like systems are equipped with LLM-empowered agents. is variable due to fluctuating task complexity and device states. The event-driven traffic in this scenario is also less sensitive to communication latency. Furthermore, the primary source of delay is the LLM inference process on robots, which may take several tens of seconds even with lightweight embedded LLMs.
@@
The third class, mission-critical control traffic, refers to time-sensitive communication flows in multi-robot systems for control purposes. Unlike time-insensitive multi-robot tasks, such as those demonstrated with the LLMind framework [25, 26], in which a high-level subtask description is assigned to a device and get locally executed under the control of the embedded AI in the robot, a time-critical control command is generated by a central controller or a master robot, and it contains specific control messages that directly regulate the action of a slave robot. Typical examples of this type of traffic include the control of high-speed motors or real-time trajectory updates for distributed robots. As the control of robotic operations requires ultra-high precision, the control message must be delivered timely to ensure the operation is accurate and updated. Even microsecond-scale delays can destabilize control loops, leading to potential system failures [10].
To support the timely distribution of mission-critical control messages from the master robot to slave robots, ROS has adopted a Publish-Subscribe (Pub/Sub) architecture built atop the Data Distribution Service (DDS) middleware. In the Pub/Sub architecture, the master node publishes control messages (e.g., velocity commands) to designated topics, typically using the Real-Time Publish-Subscribe (RTPS) protocol over TCP. Meanwhile, communication requirements are specified via Quality of Service (QoS) parameters within DDS, among which the QoS DEADLINE policy is pivotal for real-time robot controls — it specifies the maximum allowable interval between consecutive messages to enforce stringent timing requirements. For example, a 100 ms deadline requires the publisher to send an update at least every 100 ms, and the subscriber (e.g., an actuator controller that executes the command) expects to receive messages within this timing window. If this timing constraint is violated, probably because of delays in message generation or network transmission, the DDS middleware triggers events to alert the system. In such cases, the delayed command is no longer useful for execution, since the control action it represents has already expired.
@@ II-B Motivation Demonstrated via Emulation Experiments
Figure 3: A photo of the real-world system setup
@@
3) The mission-critical robot command traffic was represented by a deadline-constrained packet in ROS, with the QoS deadline request set to 100ms. That communication link was realized by the Openwifi AP and the Openwifi STA [^2] The mission-critical robot command traffic can also be simulated with commercial AP and STAs. Here we use the more complex openwifi platform to realize that traffic link to provide a fair comparison for our later experiment: our later implementation of the hybrid TDMA/CSMA system was built upon the openwifi system, simulating the pure CSMA case with the same hardware platform would provide a fair baseline for later comparisons..
@@
Traffic Type3 Description
@@
Event-driven traffic An LLMind system with the following setup: 1. Server to Client: 2 KB (Polling) 2.Client to Server: 512 KB (Historical Data)4 3. Server to Client: 5 KB (FSM Data)5
@@
    As mentioned in Section II-A, we consider a robotic traffic pattern at an overall aggregate channel load of 77.1%, wherein 99.9% of the traffic is from non time-critical applications (large-data traffic and event-driven traffic) and 0.1% of the traffic is from deadline-constraint applications.
  • 4
@@
  • 5
    FSM Data is the finite-state machine representation of decomposed tasks generated by the server to guide subsequent control script execution in LLMind system.
@@ III-A Design Principles of the Hybrid TDMA/CSMA Protocol
Third, for multi-robot cooperation, potential collisions may occur between TDMA and CSMA sections due to their different access mechanisms. CSMA’s contention‑based channel access may overlap with TDMA slots or the control slot of TDMA, causing interference under diverse traffic conditions. Therefore, an effective mechanism is required to guarantee the deterministic delivery of TDMA packets while efficiently using residual bandwidth for CSMA to manage dynamic, non-critical robot traffic.
To address these challenges, our hybrid TDMA/CSMA protocol has three key features. First, we integrate the Precision Time Protocol (PTP) for sub-microsecond synchronization across nodes to support IEEE 802.11-compatible precise alignment of TDMA slots despite robot mobility. Second, we introduce a three-section superframe that separates traffic into: (i) deterministic TDMA slots for mission-critical control traffic; (ii) a TDMA control section for transmitting dynamic slot allocation information, and (iii) a general-purpose CSMA section for non-critical robot traffic. Finally, we employ the Network Allocation Vector (NAV) mechanism, embedded in beacon frames, to protect both the TDMA section and TDMA control information by preemptively reserving the channel. This prevents interference from opportunistic CSMA accesses and guarantees the deterministic, low-latency communication essential for multi-robot cooperation.
@@ III-B1 Feature 1: PTP Synchronization Mechanism
Precise time synchronization is essential for aligning time slots in the TDMA section. We achieve a sub-microsecond clock alignment across nodes through a three-way handshake mechanism, compensating for clock offsets and propagation delays.
@@
$\tilde{s}_{i,beacon}^{l}=s_{ap,beacon}^{l}+{{d}_{i,delay}}+{{o}_{i}}$ (1)
For synchronization purposes, any one of the slave clients should respond to AP with another synchronization message (referred to as a synchronization response message in the rest of this paper) at any time during superframe section 2. We require each client accepted into the network to perform the synchronization procedure periodically. Specifically, for the synchronization process, the client embeds its transmission time, $s_{i,response}^{l}$, in the packet’s payload during the l-th superframe, representing the time the packet is sent according to the client’s local clock. The time $s_{i,response}^{l}$ is determined by the client’s local scheduling, initially based on the beacon’s arrival time, and progressively adjusted as the synchronization process continues (see Sections IV-A3 and IV-C in [21]). The server records the packet’s arrival time, $t_{i}^{l}$, and includes it in the next beacon, which is received by the client. Using these timestamps and assuming symmetric propagation delays ($d_{i}=d_{i,0}$), clients estimate:
${{d}_{i}}=\frac{1}{2}(\tilde{s}_{i,beacon}^{l}+t_{i}^{l}-s_{ap,beacon}^{l}-s_{i,response}^{l})$ (2)
${{o}_{i}}=\frac{1}{2}(\tilde{s}_{i,beacon}^{l}-t_{i}^{l}-s_{ap,beacon}^{l}+s_{i,response}^{l})$ (3)
To align packet arrivals at the server, the $i$-th clients schedules its transmission time for the $j$-th slot of the $l$-th superframe section 1 as follows:
@@
III-B2 Feature 2: Multi-Section Superframe with Flexible Time Slot Allocation